# FM4Com: Foundation Model for Scene-Adaptive Communication Strategy Optimization

### v1

@@ V Conclusion
In this paper, we proposed FM4Com, a foundation model for interactive communication strategy customization. FM4Com has been designed to comprehensively perceive both CSI and user intent. By integrating physical-layer conditions with natural-language instructions, the model first generates an intermediate chain of thought to infer user-specific preferences, and subsequently produces a customized and optimal link configuration strategy that aligns with both the current channel state and the inferred user intent, thereby enabling personalized and adaptive optimization of communication systems. During training, heuristic algorithms have been employed to collect high-quality experience samples for behavior cloning, followed by reinforcement learning fine-tuning to obtain the optimal policy. Extensive experimental evaluations have demonstrated that FM4Com significantly outperforms traditional planning-based algorithms under various challenging channel conditions, while flexibly generating optimal physical-layer strategies tailored to different user preferences and communication environments.

### v2

@@ V Conclusion
In this paper, we proposed AGENS, a foundation model for interactive communication strategy customization. AGENS has been designed to comprehensively perceive both CSI and user intent. By integrating physical-layer conditions with natural-language instructions, the model first generates an intermediate chain of thought to infer user-specific preferences, and subsequently produces a customized and optimal link configuration strategy that aligns with both the current channel state and the inferred user intent, thereby enabling personalized and adaptive optimization of communication systems. During training, heuristic algorithms have been employed to collect high-quality experience samples for SFT, followed by reinforcement learning fine-tuning to obtain the optimal policy. Extensive experimental evaluations have demonstrated that AGENS significantly outperforms traditional planning-based algorithms under various challenging channel conditions, while flexibly generating optimal physical-layer strategies tailored to different user preferences.